Card from 52 deck:

are the events being dealt a jack and being dealt a spade independent? prove
Answer #1

Ans:

Probability of choosing a jack=4/52=1/13

Probability of choosing a spade=13/52=1/4

Probability of choosing jack and spade=1/52

P(jack and spade)=P(jack)*P(spade)=(1/13)*(1/4)=1/52

Yes,independent events.

(As,P(A and B)=P(A)*P(B))